OpenStudy (anonymous):

What is the vertex of y = |x + 1| ? answer choices: (–1, 0) (0, –1) (0, 1) (1, 0)

OpenStudy (anonymous):

the vortex of y=|x| is at origin when it is shifted by amount a then general form of vortex is \[\Large y=|x-a|\] where a is the vortex so write it as \[\Large y=|x-(-1)|\] can you tell me now where the vortex is

OpenStudy (jkristia):

What is the smallest possible value of y ? |-2 + 1| = 1 |-1 + 1| = 0 |0 + 1| = 1

OpenStudy (anonymous):

is it (0, 1)?

OpenStudy (anonymous):

@jkristia @sami-21

OpenStudy (jkristia):

if x = 0, then what is y?

OpenStudy (anonymous):

-1? @jkristia

OpenStudy (jkristia):

since you equation is y = |x + 1| then if x = 1, then you have y = |1 + 1| = ?

OpenStudy (jkristia):

oops, if x = 0 then y = |0 + 1|

OpenStudy (anonymous):

I'm so confused.... idk what the answer would be

OpenStudy (anonymous):

i said... (0, –1)?

OpenStudy (anonymous):

just compare with what i mentioned early . you can see that a=-1 so vortex will be ???

OpenStudy (jkristia):

forget about the absolute value for a moment, then equation is y = x + 1 so if x = 0, then y = 0 + 1 = 1 if x = -1, then y = -1 + 1 = 0 right?

OpenStudy (anonymous):

So would it be (–1, 0)? if x= -1

OpenStudy (jkristia):

if x is -2, then the equation is y = -2 + 1 = -1, the absolute value of |-1| = 1, so the minimum y is when x = -1 x = -2, y = 1 x = -1, y = 0 x = 0, y = 1 etc...

OpenStudy (jkristia):

correct, (-1, 0)

OpenStudy (anonymous):

Oh okay... thankss @jkristia